What is the magnitude and direction of a Force with components Fx = -37kN and Fy = - 44kN?
1
Expert's answer
2021-08-26T08:10:51-0400
"F=\sqrt{37^2+44^2}=57.5\ kN\\
\theta=180+\arctan{\frac{44}{37}}=230\degree"